OpenStudy (anonymous):

log sqrt(x^3q^5) = ?

OpenStudy (anonymous):

\[\log \sqrt{x^3y^5} = ?\] expand as much possible

OpenStudy (anonymous):

\[1/2 \lg x ^{3}y ^{5}\]

OpenStudy (anonymous):

1.) Use the power rule. (For the square root/fractional exponent) 2.) Use the product rule. (Turn the multiplication into addition) 3.) Use the power rule. (For the exponents) =1/2(3logx + 5 log y)